# A Variational Approach to Planar Choreographies via Ekeland’s Principle

### v1

@@ 5 Collision avoidance via Marchal’s criterion
If a local variation $x^{\sharp}$ decreases the action in Marchal’s sense on a subinterval, then
$\mathcal{F}(x^{\sharp})=\min_{\lambda>0}\Phi_{\alpha}(\lambda;x^{\sharp})\leq\Phi_{\alpha}\!\big(\lambda^{*}(x);x^{\sharp}\big)<\Phi_{\alpha}\!\big(\lambda^{*}(x);x\big)=\mathcal{F}(x),$
where $\lambda^{*}(x)$ minimizes $\Phi_{\alpha}(\cdot;x)$. Consequently, collision exclusion also holds for minimizers of $\mathcal{F}$.
@@
Binary-collision versus regular path. For illustration, consider two equal masses undergoing a symmetric binary collision near $t=0$, with mutual distance $r(t)\sim c\,t^{2/(2+\alpha)}$ for some constant $c>0$. Then the kinetic and potential contributions satisfy $|\dot{r}|^{2}\sim t^{-2\alpha/(2+\alpha)}$ and $r^{-\alpha}\sim t^{-2\alpha/(2+\alpha)}$, so that both
$K\sim\int_{0}^{\varepsilon}t^{-2\alpha/(2+\alpha)}\,dt,\qquad V\sim\int_{0}^{\varepsilon}t^{-2\alpha/(2+\alpha)}\,dt,$
diverge as $\varepsilon\to 0$. Consequently the scale–invariant functional $\mathcal{F}=K^{\alpha/(\alpha+2)}V^{2/(\alpha+2)}$ also diverges for collision trajectories. In contrast, any nearby deformation that removes the collision keeps $r(t)\geq\delta>0$, so that $K$ and $V$ are finite and $\mathcal{F}$ attains a finite (and therefore strictly smaller) value. This heuristic computation illustrates the mechanism behind Marchal’s criterion: collision paths cannot minimize $\mathcal{F}$ since any local regular perturbation yields a smaller action value.
@@ 6 Regularity of the critical point
The minimizer $x_{\ast}$ is collision–free; hence $\nabla U$ and all its derivatives are smooth and bounded along the trajectory. The weak equation $m\ddot{x}_{\ast}=\rho\nabla U(x_{\ast})$ then yields, by an iterative regularity argument, that $x_{\ast}\in C^{\infty}([0,2\pi])$ (and even real–analytic if $U$ is analytic away from collisions).
@@
Since $x_{\ast}\in H^{1}(\mathbb{T};\mathbb{R}^{2N})$ and satisfies
$m\int_{0}^{2\pi}\dot{x}_{\ast}\cdot\dot{\eta}\,dt=\rho\int_{0}^{2\pi}\nabla U(x_{\ast})\cdot\eta\,dt\qquad\forall\,\eta\in C^{\infty}_{\mathrm{per}},$
we have in the distributional sense $m\,\ddot{x}_{\ast}=\rho\,\nabla U(x_{\ast})$. Because the curve avoids the singular set, the image $K=x_{\ast}([0,2\pi])$ is a compact set on which $\nabla U$ is $C^{\infty}$ and bounded. Thus $\nabla U(x_{\ast})\in L^{\infty}(0,2\pi)$, so $\ddot{x}_{\ast}\in L^{\infty}$ and $x_{\ast}\in W^{2,\infty}(0,2\pi)\hookrightarrow C^{1}$.
From here a standard bootstrap applies: if $x_{\ast}\in C^{r}$ and $U\in C^{r+1}$ on $K$, then $\nabla U(x_{\ast})\in C^{r}$ and hence $x_{\ast}\in C^{r+2}$. Since $U$ is $C^{\infty}$ off collisions, we obtain inductively $x_{\ast}\in C^{\infty}$. If, in addition, $U$ is analytic, the differential equation is analytic and $x_{\ast}$ is analytic as well. ∎

### v2

@@ 3.1 From the scale–invariant functional to Newton’s equation
Remark 3.5 (Equivalence of critical points of $\mathcal{A}$ and $\mathcal{F}$).
Although the functionals $\mathcal{A}$ and $\mathcal{F}$ are not identical, they have the same critical points up to time–reparametrization. Indeed, every critical point of $\mathcal{A}$ is mapped to a critical point of $\mathcal{F}$ by choosing the minimising scale $\lambda>0$, and conversely every critical point of $\mathcal{F}$ yields, after the rescaling $s=\lambda t$, a solution of Newton’s equation. Thus $\mathcal{F}$ is a scale–invariant reformulation of the classical action functional $\mathcal{A}$, adapted to the variational treatment of choreographies.
@@ 4 Ekeland’s principle: statement and application to $\mathcal{F}$
Since $\mathcal{M}=\{K=1\}$ is a $C^{1}$ codimension–one submanifold of the Hilbert space $X=H^{1}(\mathbb{T};(\mathbb{R}^{2})^{N})$, its tangent space at $x\in\mathcal{M}$ is
$T_{x}\mathcal{M}=\{\eta\in X:\ DK(x)[\eta]=0\}.$
By the Riesz representation in $X$, the differential $DK(x)$ can be written as
$DK(x)[\eta]=\langle\nabla K(x),\eta\rangle_{H^{1}},$
for a unique vector $\nabla K(x)\in X$, so that
$T_{x}\mathcal{M}=\{\eta\in X:\ \langle\nabla K(x),\eta\rangle_{H^{1}}=0\}=\big(\operatorname{span}\{\nabla K(x)\}\big)^{\perp}.$
Thus we have the orthogonal decomposition
$X=T_{x}\mathcal{M}\oplus\operatorname{span}\{\nabla K(x)\}.$
For each $x\in\mathcal{M}$ and each $v\in X$, there exists a unique scalar $\mu\in\mathbb{R}$ such that
$v=v_{T}+\mu\,\nabla K(x),\qquad v_{T}\in T_{x}\mathcal{M}.$
Applied to $v=\nabla f(x)$, this gives the decomposition
$\nabla f(x)=\nabla_{\mathcal{M}}f(x)+\mu(x)\,\nabla K(x),$
where $\nabla_{\mathcal{M}}f(x)\in T_{x}\mathcal{M}$ is the tangential gradient of $f$ along $\mathcal{M}$.
In particular, for the sequence $\{x_{n}\}\subset\mathcal{M}$ given by Ekeland’s principle we may write
$\nabla f(x_{n})+\mu_{n}\,\nabla K(x_{n})=\nabla_{\mathcal{M}}f(x_{n}).$
Since Ekeland yields $\|\nabla_{\mathcal{M}}f(x_{n})\|\leq 1/n$, we obtain
$\|\nabla f(x_{n})+\mu_{n}\,\nabla K(x_{n})\|=\|\nabla_{\mathcal{M}}f(x_{n})\|\longrightarrow 0,$
which is precisely a Palais–Smale type condition for the restriction of $f$ to $\mathcal{M}$.
@@ 5 Collision avoidance via Marchal’s criterion
For each $x$ let $\lambda^{*}(x)>0$ be the (unique) minimizer of $\Phi_{\alpha}(\cdot;x)$, and set
$y:=x_{\lambda^{*}(x)}.$
Then, by the scaling laws for $K$ and $V$,
$\mathcal{F}(x)=\min_{\lambda>0}\Phi_{\alpha}(\lambda;x)=\Phi_{\alpha}\big(\lambda^{*}(x);x\big)=\mathcal{A}(y).$
Suppose now that $y$ has a collision. By Marchal’s construction there exists a local variation $y^{\sharp}$ of $y$ (preserving the endpoints and global symmetries) such that
$\mathcal{A}(y^{\sharp})<\mathcal{A}(y)=\Phi_{\alpha}\big(\lambda^{*}(x);x\big)=\mathcal{F}(x).$
Define $x^{\sharp}$ by $y^{\sharp}=(x^{\sharp})_{\lambda^{*}(x)}$. We claim that $\mathcal{F}(x^{\sharp})<\mathcal{F}(x)$, and we justify each step in the chain
$\mathcal{F}(x^{\sharp})=\min_{\lambda>0}\Phi_{\alpha}(\lambda;x^{\sharp})\leq\Phi_{\alpha}\big(\lambda^{*}(x);x^{\sharp}\big)=\mathcal{A}(y^{\sharp})<\mathcal{A}(y)=\Phi_{\alpha}\big(\lambda^{*}(x);x\big)=\mathcal{F}(x).$
(i) Definition of $\mathcal{F}(x^{\sharp})$. By definition,
$\mathcal{F}(x^{\sharp})=\min_{\lambda>0}\Phi_{\alpha}(\lambda;x^{\sharp}).$
(ii) Evaluation at a fixed scale. For any fixed $\lambda>0$ one has
$\min_{\lambda>0}\Phi_{\alpha}(\lambda;x^{\sharp})\leq\Phi_{\alpha}(\lambda;x^{\sharp}).$
In particular, taking $\lambda=\lambda^{*}(x)$ gives
$\min_{\lambda>0}\Phi_{\alpha}(\lambda;x^{\sharp})\leq\Phi_{\alpha}\big(\lambda^{*}(x);x^{\sharp}\big).$
Here equality does not hold in general, since $\lambda^{*}(x)$ minimizes $\Phi_{\alpha}(\cdot;x)$ but need not minimize $\Phi_{\alpha}(\cdot;x^{\sharp})$.
(iii) Identification with the action of $y^{\sharp}$. By definition of $y^{\sharp}=(x^{\sharp})_{\lambda^{*}(x)}$ and of $\Phi_{\alpha}$,
$\Phi_{\alpha}\big(\lambda^{*}(x);x^{\sharp}\big)=\mathcal{A}(y^{\sharp}).$
(iv) Marchal’s strict inequality. Marchal’s deformation gives
$\mathcal{A}(y^{\sharp})<\mathcal{A}(y).$
(v) Identification with $\mathcal{F}(x)$. Since $\lambda^{*}(x)$ minimizes $\Phi_{\alpha}(\cdot;x)$, we have
$\mathcal{A}(y)=\Phi_{\alpha}\big(\lambda^{*}(x);x\big)=\mathcal{F}(x).$
Combining (i)–(v) yields $\mathcal{F}(x^{\sharp})<\mathcal{F}(x)$. Hence any curve with an interior collision can be locally modified to produce a strictly smaller value of $\mathcal{F}$, so collision exclusion also holds for minimizers of the scale–invariant functional $\mathcal{F}$.
@@
Binary collision versus regular path. For illustration, consider two equal masses undergoing a symmetric binary collision near $t=0$, with mutual distance $r(t)\sim c\,t^{2/(2+\alpha)}$ for some constant $c>0$. Then $|\dot{r}(t)|^{2}$ and $r(t)^{-\alpha}$ have the same local behaviour $t^{-2\alpha/(2+\alpha)}$, so that their contributions to the action over $(0,\varepsilon)$ scale like
$K_{\varepsilon}\sim\int_{0}^{\varepsilon}t^{-2\alpha/(2+\alpha)}\,dt,\qquad V_{\varepsilon}\sim\int_{0}^{\varepsilon}t^{-2\alpha/(2+\alpha)}\,dt\sim C\,\varepsilon^{(2-\alpha)/(2+\alpha)},$
which remains finite for $0<\alpha<2$. Thus collision arcs have finite action, and a direct divergence argument cannot be used to discard them. Instead, Marchal’s construction shows that one can perform a local deformation near the collision —preserving the endpoints and the global symmetries— which strictly lowers the action $\mathcal{A}$. By Remark 5.2, this implies a strict decrease of the scale–invariant functional $\mathcal{F}$ as well, so collision paths cannot be minimizers of $\mathcal{F}$.
@@ 6 Regularity of the critical point
The minimizer $x_{\ast}$ is collision–free. Therefore $U$ and all its derivatives are smooth and bounded along the trajectory. The weak Newton equation
$m\,\ddot{x}_{\ast}=\nabla U(x_{\ast})$
then implies, via a standard iterative bootstrap argument, that $x_{\ast}\in C^{\infty}([0,2\pi])$; moreover, if $U$ is analytic away from collisions, then $x_{\ast}$ is real–analytic in time.
@@
After the time–rescaling $s=\sqrt{\rho}\,t$ performed in the previous section, the minimizer satisfies Newton’s equation in the classical form
$m\,\ddot{x}_{\ast}=\nabla U(x_{\ast}),$
where the dot denotes differentiation with respect to the rescaled time $s$. We keep the notation $t$ for simplicity.
Step 1: Weak formulation.
Since $x_{\ast}\in H^{1}(\mathbb{T};\mathbb{R}^{2N})$ minimizes the action under the symmetry constraints, for every test function $\eta\in C^{\infty}_{\mathrm{per}}([0,2\pi];\mathbb{R}^{2N})$ we have
$m\int_{0}^{2\pi}\dot{x}_{\ast}(t)\cdot\dot{\eta}(t)\,dt=\int_{0}^{2\pi}\nabla U(x_{\ast}(t))\cdot\eta(t)\,dt.$ (11)
Integrating by parts on the left (boundary terms vanish by periodicity), we obtain
$m\int_{0}^{2\pi}\ddot{x}_{\ast}(t)\cdot\eta(t)\,dt=\int_{0}^{2\pi}\nabla U(x_{\ast}(t))\cdot\eta(t)\,dt\qquad\forall\,\eta.$
Thus, in the sense of distributions,
$m\,\ddot{x}_{\ast}=\nabla U(x_{\ast}).$
Step 2: $L^{\infty}$ regularity and $C^{1}$ regularity.
Because $x_{\ast}$ has no collisions, its image
$K:=x_{\ast}([0,2\pi])\subset(\mathbb{R}^{2})^{N}\setminus\Delta$
is compact and lies at positive distance from the singular set $\Delta$. Hence
$\nabla U|_{K}\in C^{\infty}(K),\qquad\|\nabla U\|_{L^{\infty}(K)}<\infty.$
Therefore the right-hand side $\nabla U(x_{\ast}(t))$ belongs to $L^{\infty}(0,2\pi)$, so
$\ddot{x}_{\ast}\in L^{\infty}(0,2\pi),\qquad x_{\ast}\in W^{2,\infty}(0,2\pi).$
Since in one dimension
$W^{2,\infty}(0,2\pi)\hookrightarrow C^{1,1}([0,2\pi]),$
we obtain
$x_{\ast}\in C^{1}([0,2\pi])\quad\text{and}\quad\dot{x}_{\ast}\text{ is Lipschitz}.$
Step 3: Start of the bootstrap.
Having $x_{\ast}\in C^{1}$ and $\nabla U\in C^{\infty}$ on $K$, the chain rule implies
$t\longmapsto\nabla U(x_{\ast}(t))\in C^{1}([0,2\pi]).$
Then the ODE
$m\,\ddot{x}_{\ast}=\nabla U(x_{\ast})$
shows that $\ddot{x}_{\ast}\in C^{1}$, hence
$x_{\ast}\in C^{3}([0,2\pi]).$
Step 4: Inductive step.
Assume $x_{\ast}\in C^{r}$ for some $r\geq 1$. Because $U\in C^{\infty}$ on $K$, all derivatives $\partial^{\beta}\nabla U$ are bounded on $K$. Composing with $x_{\ast}\in C^{r}$ yields
$\nabla U(x_{\ast})\in C^{r}.$
Thus $\ddot{x}_{\ast}\in C^{r}$, and integrating twice gives
$x_{\ast}\in C^{r+2}.$
Step 5: Conclusion of the bootstrap.
Starting from $x_{\ast}\in C^{1}$ we obtain successively
$C^{1}\longrightarrow C^{3}\longrightarrow C^{5}\longrightarrow\cdots$
and therefore
$x_{\ast}\in C^{\infty}([0,2\pi];\mathbb{R}^{2N}).$
Step 6: Analyticity.
If $U$ is real–analytic on $(\mathbb{R}^{2})^{N}\setminus\Delta$, then $\nabla U$ is analytic, and the ODE
$m\,\ddot{x}_{\ast}=\nabla U(x_{\ast})$
is analytic on the compact collision–free set $K$. By the classical Cauchy–Kowalevski theorem for analytic ODEs, any $C^{2}$ solution whose image stays in $K$ is analytic in $t$. Since we have already shown $x_{\ast}\in C^{\infty}$ and $x_{\ast}([0,2\pi])\subset K$, it follows that $x_{\ast}$ is analytic.